# 4.2cm4.2cm3.6cm3cm

### v1

@@
Abstract: In this paper, generalized Loop Heisenberg-Virasoro algebras are introduced. Firstly, we determine the derivations on generalized Loop Heisenberg-Virasoro algebras. Then we show that all $2$-local derivations are derivations. Furthermore, we determine the biderivations on generalized Loop Heisenberg-Virasoro algebras are inner biderivations and give their applications. Finally, the automorphism groups on generalized Loop Heisenberg-Virasoro algebras are presented.
@@ Introduction
As is well known, structure theory is essential in the research of Lie algebras. Likewise, the derivation algebra and automorphism group on the generalized Loop Virasoro algebra, the generalized Loop Schrödinger–Virasoro algebras and the generalized Heisenberg-Virasoro algebra, etc. have been determined respectively (see [6, 7, 9, 10, 19]). $2$-local derivations on the generalized Witt algebras are also obtained in [5]. The author of [20] proved that all 2-local derivations on the Witt algebra as well as on the positive Witt algebra are derivations, and give an example of infinite-dimensional Lie algebra with a 2-local derivation which is not a derivation. The author of [3] determined biderivations on $m_{0},\,L_{1},\,m_{2}$ which are three Lie algebras of maximal class with one-dimensional homogeneous components, and characterized their local and 2-local derivations. All biderivations on the twisted Heisenberg–Virasoro algebra, the Schrödinger-Virasoro algebra and the W-algebras, etc. have been given in [11, 12, 13, 18].
Likewise, representation theory is also very important in the research of Lie algebras. For example, the author of [2] described the structure of the irreducible highest weight modules for the twisted Heisenberg-Virasoro Lie algebra at level zero, sufficient and necessary conditions are given for these non-weight modules on the mirror Heisenberg-Virasoro algebra to be irreducibie (see [14]), the author of [15, 16, 17] also investigated representations on the mirror Heisenberg-Virasoro algebra.
@@
Inspired by [8], we introduced the generalized Loop Heisenberg-Virasoro algebra. When $\Gamma=\mathbb{Z}$, the generalized Loop Heisenberg-Virasoro algebra in the paper is the Loop Heisenberg-Virasoro algebra [8]. The aim is to determine the derivations, $2$-local derivations, biderivations and automorphism groups of generalized Loop Heisenberg-Virasoro algebra by use of the methods for derivations, $2$-local derivations, biderivations, and automorphism groups of respective Lie algebras in [7, 9, 12, 20]. The structure of this paper is as follows: Firstly, based on its relevant properties, the derivations on generalized Loop Heisenberg-Virasoro algebra are determined. On this basis, it is proved that all $2$-local derivations of this algebra are derivations. Secondly, based on the fact that it is a perfect Lie algebra, we conclude that all biderivations are inner biderivations. Then we prove the necessary and sufficient conditions for its commutating linear maps and any commutative $post$-Lie algebra structure on this algebra is trivial. Finally, we define five types of automorphism groups on it and prove that any automorphism group of the algebra can be represented the direct product of the five types of automorphisms.
@@ 1. Preliminaries
$[L_{\alpha,i},L_{\beta,j}]$ $=$ $(\alpha-\beta)L_{\alpha+\beta,i+j},$
$[L_{\alpha,i},H_{\beta,j}]$ $=$ $-\beta H_{\alpha+\beta,i+j},$
$[H_{\alpha,i},H_{\beta,j}]$ $=$ $0$
In the sequel, generalized Loop Heisenberg-Virasoro algebra denoted by $\mathcal{L}(\Gamma)$. It is worth noting that Loop Heisenberg-Virasoro algebra in [8] is exactly $\mathcal{L}(\Gamma)$ when $\mathbb{Z}=\Gamma$. Let $\mathcal{LV}$ be subalgebra of $\mathcal{L}(\Gamma)$ spanned by $\{L_{\alpha,i}\,|\,\alpha\in\Gamma,\,i\in\mathbb{Z}\}$. Then $\mathcal{LV}$ is actually isomorphic to the centerless generalized Loop Virasoro algebra (see [7]). In addition, $\mathcal{L}(\Gamma)$ contains the generalized Heisenberg-Virasoro algebra
@@
$\mathcal{D}([x,y])=x\cdot\mathcal{D}(y)+y\cdot\mathcal{D}(x),\mbox{for all}\,x,y\in L.$
A linear map $\mathrm{ad}:x\mapsto x\cdot v$ is called the inner derivation. Otherwise others are called outer derivations.
@@
$\mathcal{L}(\Gamma)_{\mu}=\left\{x\in\mathcal{L}(\Gamma)\,|\,[L_{0,0},x]=-\mu x\right\}=\left\{L_{\mu,j},H_{\mu,j}\,|\,j\in\mathbb{Z}\right\}.$
@@ 2. Derivations on $\mathcal{L}(\Gamma)$
$\mathrm{Der}\mathcal{L}(\Gamma)=(\mathrm{Der}\mathcal{L}(\Gamma))_{0}\bigoplus\mathrm{ad}\mathcal{L}(\Gamma)$.
@@
(i) For all $\mathcal{D}\in\mathrm{Der}\mathcal{L}(\Gamma),x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha}$, assume that $\mathcal{D}(x_{\alpha})=\sum_{\beta\in\Gamma}y_{\beta}$. We define $\mathcal{D}_{\gamma}(x_{\alpha})=y_{\alpha+\gamma}$, where $\mathcal{D}_{\gamma}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$. It is easy to prove that
$\mathcal{D}_{\gamma}([x_{\alpha},x_{\beta}])=[\mathcal{D}_{\gamma}(x_{\alpha}),x_{\beta}]+[x_{\alpha},\mathcal{D}_{\gamma}(x_{\beta})],\,\forall x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha},x_{\beta}\in\mathcal{L}(\Gamma)_{\beta}.$
Let $x=\sum_{\alpha\in\Gamma}x_{\alpha}$, where $x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha}$. We can conclude that
$\mathcal{D}(x)=\sum_{\alpha\in\Gamma}\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x_{\alpha})=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x).$
That is for any derivation $\mathcal{D}\in\mathcal{L}(\Gamma)$, one has
$\mathcal{D}=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma},\,\mathcal{D}_{\gamma}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}.$
such that for all $x\in\mathcal{L}(\Gamma)$, only finitely $\mathcal{D}_{\gamma}(x)\neq 0$, and $\mathcal{D}(x)=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x)$.
(ii) For all $x\in\mathcal{L}(\Gamma)$, one hand it has
$\mathcal{D}([L_{0,0},x])$ $=$ $[\mathcal{D}(L_{0,0}),x]+[L_{0,0},\mathcal{D}(x)]$
$=$ $[\mathcal{D}(L_{0,0}),x]+[L_{0,0},y_{\mu+\gamma}]$
$=$ $[\mathcal{D}(L_{0,0}),x]+{-(\mu+\gamma)}\mathcal{D}(x)$,
other hand
$\mathcal{D}([L_{0,0},x])$ $=$ $\mathcal{D}(-\mu x)$
$=$ $-\mu\mathcal{D}(x)$,
@@
$\mathcal{D}(x)$ $=$ $[\frac{1}{\gamma}\mathcal{D}(L_{0,0}),x]$
$=$ ${\rm{ad}}_{\gamma^{-1}\mathcal{D}(L_{0,0})}(x)$.
Thus $\mathcal{D}=\mathrm{ad}_{\gamma^{-1}\mathcal{D}(L_{0,0})}$ is an inner derivation. When $\gamma\neq 0$, for all $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$ is an inner derivation.
(iii) For all $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}\cap\mathrm{ad}\mathcal{L}(\Gamma)$. On the one hand, there exists an $X\in\mathcal{L}(\Gamma)$ such that $\mathcal{D}(Y_{\alpha,i})=[X,Y_{\alpha,i}]$, where $Y\in\mathcal{L}(\Gamma),\alpha\in\Gamma,\,i\in\mathbb{Z}$.
On the other hand, by $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$, then for $Z_{\gamma,k}\in\mathcal{L}(\Gamma)_{\gamma}$, where $\gamma\in\Gamma,k\in\mathbb{Z}$, we have that $\mathcal{D}(Z_{\gamma,k})=[X,Z_{\gamma,k}]\in\mathcal{L}(\Gamma)_{\gamma}$.
By $X=\sum_{\begin{subarray}{c}\beta\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}X_{\beta,j}$, where $X_{\beta,j}\in\mathcal{L}(\Gamma)_{\beta}$, we have
$[X,Z_{\gamma,k}]=\sum_{\begin{subarray}{c}\beta\in\Gamma\\ j\in\mathbb{Z}\end{subarray}}[X_{\beta,j},Z_{\gamma,k}]\in\mathcal{L}(\Gamma)_{\gamma}.$
Thus $[X_{\beta,j},Z_{\gamma,k}]=0,\,\forall\beta\in\Gamma,\,j\in\mathbb{Z}$, that is $\mathcal{D}=0$. ∎
By Lemma 2, if we want to determine $\mathrm{Der}\mathcal{L}(\Gamma)$, only needs to define $(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. For all $\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),g\in\mathbf{g}(\Gamma),b\in\mathbb{F}[t^{1},t^{-1}],\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}$ and
$\mathbf{g}(\Gamma)=\left\{g:\Gamma\rightarrow\mathbb{F}[t^{1},t^{-1}]\,|\,(\alpha-\beta)g_{\alpha+\beta}=\alpha g_{\alpha}-\beta g_{\beta},\,\forall\alpha,\beta\in\Gamma\right\},$
where $\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}$ is a derivation algebra of $\mathbb{F}[t^{1},t^{-1}]$. We brief note $\mathbf{g}(\Gamma)$ by $\mathbf{g}$ in the sequel. The following linear maps $D_{\phi},\,D_{g},\,D_{b},\,D^{\rho}$ on $\mathcal{L}(\Gamma)$ are given by respectively:
$D_{\phi}(X_{\mu,i})$ $=$ $\phi(\mu)X_{\mu,i}$,
$D_{g}(L_{\alpha,i})$ $=$ $g_{\alpha}H_{\alpha,i},\quad D_{g}(H_{\alpha,i})=0$,
$D_{b}(L_{\alpha,i})$ $=$ $0,\quad D_{b}(H_{\alpha,i})=bH_{\alpha,i}$,
$D^{\rho}(X_{\mu,i})$ $=$ $X_{\mu}\rho(t^{i}),\,\forall X\in\{L,H\},\,i\in\mathbb{Z},\,\mu,\alpha\in\Gamma$.
@@
$\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])}\,:=\,\left\{D_{\phi}\,|\,\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}])\right\}$,
@@
$\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]}\,:=\,\left\{D_{b}\,|\,b\in\mathbb{F}[t^{1},t^{-1}]\right\}$,
$\mathcal{D}_{\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}}\,:=\,\{D^{\rho}\,|\,\rho\in\mathbb{F}[t^{1},t^{-1}]\frac{d}{dt}\}$.
@@
Let $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. Then there exists $f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]),\,g\in\mathbf{g}$ such that
$D(L_{\alpha,i})=f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i},\,\alpha\in\Gamma,i\in\mathbb{Z}.$
@@
Let $D(L_{\alpha,i})=f_{\alpha,i}L_{\alpha,i}+g_{\alpha,i}H_{\alpha,i},\,f_{\alpha,i},g_{\alpha,i}\in\mathbb{F}[t^{1},t^{-1}]$. We have
$D([L_{\alpha,i},L_{\beta,j}])$ $=$ $D({(\alpha-\beta)L_{\alpha+\beta,i+j}})$
$=$ $(\alpha-\beta)D(L_{\alpha+\beta,i+j})$
$=$ $(\alpha-\beta)(f_{\alpha+\beta,i+j}L_{\alpha+\beta,i+j}+g_{\alpha+\beta,i+j}H_{\alpha+\beta,i+j})$.
@@
$D([L_{\alpha,i},L_{\beta,j}])$ $=$ $[D(L_{\alpha,i}),L_{\beta,j}]+[L_{\alpha,i},D(L_{\beta,j})]$
$=$ $[f_{\alpha,i}L_{\alpha,i}+g_{\alpha,i}H_{\alpha,i},L_{\beta,j}]+[L_{\alpha,i},f_{\beta,j}L_{\beta,j}+g_{\beta,j}H_{\beta,j}]$
$=$ $(\alpha-\beta)f_{\alpha,i}L_{\alpha+\beta,i+j}+\alpha g_{\alpha,i}H_{\alpha+\beta,i+j}+(\alpha-\beta)f_{\beta,j}L_{\alpha+\beta,i+j}$
$-\beta g_{\beta,j}H_{\alpha+\beta,i+j}$.
We obtain
$f_{\alpha+\beta,i+j}=f_{\alpha,i}+f_{\beta,j},\,\alpha\neq\beta$
$(\alpha-\beta)g_{\alpha+\beta,i+j}=\alpha g_{\alpha,i}-\beta g_{\beta,j},\,\forall\alpha,\beta\in\Gamma,\,i,j\in\mathbb{Z}$.
@@
Thus $f_{0,0}=0$, Let $\beta=-\alpha\neq 0,i=j=0$. Then
$0=f_{0,0}=f_{\alpha,0}+f_{-\alpha,0},\,\forall\alpha\neq 0$
Let $0\neq\eta\in\Gamma$. Satisfying $\eta\neq\alpha,\,-\eta\neq\beta,\alpha+\eta\neq\beta-\alpha$, then
@@
That is
$f_{\alpha+\beta,i+j}=f_{\alpha,i}+f_{\beta,j},\,\forall\alpha,\beta\in\Gamma,i,j\in\mathbb{Z}.$
Then $f_{\alpha,i}=f_{\alpha,0}+f_{0,i}$. Let $\rho=tf_{0,1}\frac{d}{dt}$. We have
@@
Replace $D$ by $D-D^{\rho}$. Let $D(L_{0,1})=0$. Then $f_{0,1}=0$ and $f_{0,i}=if_{0,1}=0$, thus
$f_{\alpha,i}=f_{\alpha,0}+f_{0,i}=f_{\alpha,0},\,\forall\alpha\in\Gamma,i\in\mathbb{Z}.$
For convenience, denote $f_{\alpha,i}$ as $f_{\alpha}$, then
$f_{\alpha+\beta}=f_{\alpha}+f_{\beta},\,f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t^{1},t^{-1}]).$
Let $\alpha=0,j=0$. We have
@@
Let $\alpha=-\beta\neq 0$. We have
$2\beta g_{0,i+j}$ $=$ $\beta g_{\beta,j}+\beta g_{-\beta,i}$
$=$ $\beta g_{\beta,0}+\beta g_{-\beta,0}$
$=$ $2\beta g_{0,0}$
$=$ 00.
Thus
$g_{\beta,i}=g_{\beta,0}-g_{0,i}=g_{\beta,0},\,\forall\beta\in\Gamma,i\in\mathbb{Z}.$
For convenience, denote $g_{\beta}$ as $g_{\beta,i}$, then
$(\alpha-\beta)g_{\alpha+\beta}=\alpha g_{\alpha}-\beta g_{\beta},\,\forall g\in\mathbf{g}(\Gamma).$
@@
Let $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. Then there exists the $b\in\mathbb{F}[t^{1},t^{-1}]$ such that
$D(H_{\alpha,i})=(f_{\alpha}+b)H_{\alpha,i},\,\alpha\in\Gamma,i\in\mathbb{Z}.$

### v2

@@
Abstract: In this paper, the generalized loop Heisenberg-Virasoro algebra is introduced. Firstly, we determine the derivations on the generalized loop Heisenberg-Virasoro algebra. Then we show that all $2$-local derivations are derivations for the generalized loop Heisenberg-Virasoro algebra. Furthermore, we prove the biderivations on the generalized loop Heisenberg-Virasoro algebra are inner biderivations and give their applications. Finally, the automorphism groups on the generalized loop Heisenberg-Virasoro algebra are presented.
@@ Introduction
As is well known, structure theory is essential in the research of Lie algebras. Likewise, the authors in [16, 19, 6, 13, 14] determined the derivation algebra and automorphism group on the generalized loop Virasoro algebra, the generalized loop Schrödinger–Virasoro algebras and the generalized Heisenberg-Virasoro algebra, etc. The authors in [2] obtain $2$-local derivations on the generalized Witt algebras. The authors in [3] proved that all 2-local derivations on the Witt algebra as well as on the positive Witt algebra are derivations, and provided an example of infinite-dimensional Lie algebra with a 2-local derivation which is not a derivation. The authors in [21] determined biderivations on the three Lie algebras of maximal class with one-dimensional homogeneous components, and characterized their local and 2-local derivations. The authors in [18, 20, 17, 11] gave the biderivations on the twisted Heisenberg–Virasoro algebra, the Schrödinger-Virasoro algebra and the W-algebras. The authors in [4] described all biderivations of a semisimple Lie algebra. The authors in [5] described the structure of the irreducible highest weight modules for the twisted Heisenberg-Virasoro Lie algebra at level zero. The authors in [7] constructed two classes of non-weight modules over the twisted Heisenberg-Virasoro algebra and presented the necessary and sufficient conditions under which modules in these two classes are irreducible and isomorphic. The authors in [9] gave sufficient and necessary conditions for these non-weight modules on the mirror Heisenberg-Virasoro algebra to be irreducibie. The authors in [12] studied irreducible modules over the mirror Heisenberg-Virasoro algebra and determined irreducible modules with finite-dimensional weight spaces are either an irreducible highest or an irreducible lowest weight module, or an irreducible module of the intermediate series. The authors in [10] determined the necessary and sufficient conditions for the tensor products of irreducible highest weight modules and irreducible modules of intermediate series over the mirror Heisenberg-Virasoro algebra to be irreducible by using “shifting technique”.
@@
In this paper, inspired by [8], we introduce the generalized loop Heisenberg-Virasoro algebra. The aim is to determine the derivations, $2$-local derivations, biderivations and automorphism groups of the generalized loop Heisenberg-Virasoro algebra by use of the analogous methods with those of respective Lie algebras in [19, 6, 20, 3, 13]. The structure of this paper is as follows: Firstly, based on its relevant properties, the derivations on the generalized loop Heisenberg-Virasoro algebra are determined. On this basis, it is proved that all $2$-local derivations of this algebra are derivations. Secondly, based on the fact that it is a perfect Lie algebra, we conclude that all biderivations are inner biderivations. Then we prove the necessary and sufficient conditions for its commutating linear maps and any commutative postpost-Lie algebra structure on this algebra is trivial. Finally, we define five types of automorphism groups of the algebra and prove that any automorphism group on it can be represented the direct product of the five types of automorphisms.
@@ 1. Preliminaries
$L_{\alpha,i},L_{\beta,j}]$ == $(\alpha-\beta)L_{\alpha+\beta,i+j},$
$L_{\alpha,i},H_{\beta,j}]$ == $\beta H_{\alpha+\beta,i+j},$
$H_{\alpha,i},H_{\beta,j}]$ == $0.$
In the sequel, denoted by $\mathcal{L}(\Gamma)$ generalized loop Heisenberg-Virasoro algebra. It is worth noting that loop Heisenberg-Virasoro algebra in [8] is exactly $\mathcal{L}(\Gamma)$ when $\mathbb{Z}=\Gamma$. Let $\mathcal{LV}$ be subalgebra of $\mathcal{L}(\Gamma)$ spanned by $\{L_{\alpha,i}\,|\,\alpha\in\Gamma,\,i\in\mathbb{Z}\}$. Then $\mathcal{LV}$ is actually isomorphic to the centerless generalized loop Virasoro algebra (see [19]). In addition, $\mathcal{L}(\Gamma)$ contains the generalized Heisenberg-Virasoro algebra (see [13])
@@
$\mathcal{D}([x,y])=x\cdot\mathcal{D}(y)+y\cdot\mathcal{D}(x)\,\mbox{for all}\,x,y\in L.$
A linear map $\mathrm{ad}:x\mapsto x\cdot v$ for all $x\in L$ and all $v\in V$ is called the inner derivation. Otherwise others are called outer derivations.
@@
$\mathcal{L}(\Gamma)_{\mu}=\left\{x\in\mathcal{L}(\Gamma)\,|\,[L_{0,0},x]=-\mu x\right\}=\mathrm{span}_{\mathbb{F}}\left\{L_{\mu,j},H_{\mu,j}\,|\,j\in\mathbb{Z}\right\}.$
@@ 2. Derivations on $\mathcal{L}(\Gamma)$
$\mathrm{Der}\mathcal{L}(\Gamma)=(\mathrm{Der}\mathcal{L}(\Gamma))_{0}+\mathrm{ad}\mathcal{L}(\Gamma)$.
@@
(i) First we show that $\mathrm{Der}\mathcal{L}(\Gamma)=\sum_{\gamma\in\Gamma}(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}.$
For all $\mathcal{D}\in\mathrm{Der}\mathcal{L}(\Gamma)$ and $x=\sum_{\alpha\in\Gamma}x_{\alpha}$, where $x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha}$, assume that $\mathcal{D}(x_{\alpha})=\sum_{\beta\in\Gamma}y_{\beta}$. We define $\mathcal{D}_{\gamma}(x_{\alpha})=y_{\alpha+\gamma}$. It is easy to prove that
$\mathcal{D}_{\gamma}([x_{\alpha},x_{\beta}])=[\mathcal{D}_{\gamma}(x_{\alpha}),x_{\beta}]+[x_{\alpha},\mathcal{D}_{\gamma}(x_{\beta})],\,\mbox{for all}\,x_{\alpha}\in\mathcal{L}(\Gamma)_{\alpha}\mbox{and all}\,x_{\beta}\in\mathcal{L}(\Gamma)_{\beta}.$
Then $\mathcal{D}_{\gamma}$ is a derivation on $\mathcal{L}(\Gamma)$. We can conclude that
$\mathcal{D}(x)=\sum_{\alpha\in\Gamma}\mathcal{D}(x_{\alpha})=\sum_{\alpha\in\Gamma}\sum_{\beta\in\Gamma}y_{\beta}=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}(x).$
Then $\mathcal{D}=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}$. That is for any derivation $\mathcal{D}\in\mathrm{Der}\mathcal{L}(\Gamma)$, one has $\mathcal{D}=\sum_{\gamma\in\Gamma}\mathcal{D}_{\gamma}$, where $\mathcal{D}_{\gamma}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$, such that for all $x\in\mathcal{L}(\Gamma)$, only finitely $\mathcal{D}_{\gamma}(x)\neq 0$.
(ii) Then we show that for all $\gamma\neq 0$, $\mathcal{D}\in(\mathrm{Der}\mathcal{L}(\Gamma))_{\gamma}$ is an inner derivation.
For all $x\in\mathcal{L}(\Gamma)_{\mu}$, on the one hand, it has
$\mathcal{D}([L_{0,0},x])=[\mathcal{D}(L_{0,0}),x]{-(\mu+\gamma)}\mathcal{D}(x),$
on the other hand,
$\mathcal{D}([L_{0,0},x])$ == $\mathcal{D}(-\mu x)$
== $\mu\mathcal{D}(x)$,
@@
$\mathcal{D}(x)$ == $\frac{1}{\gamma}\mathcal{D}(L_{0,0}),x]$
== ${\rm{ad}}_{\gamma^{-1}\mathcal{D}(L_{0,0})}(x)$.
Thus for all $\gamma\neq 0$, $\mathcal{D}=\mathrm{ad}_{\gamma^{-1}\mathcal{D}(L_{0,0})}$ is an inner derivation.
By Lemma 2, if we determine $\mathrm{Der}\mathcal{L}(\Gamma)$, only needs to determine $(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. For a map $f$ from a set $A$ to another set $B$, let $f_{a}$ denote the image of $a\in A$ under $f$. For all
$\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}]),g\in\mathbf{g}(\Gamma),b\in\mathbb{F}[t,t^{-1}],\rho\in\mathbb{F}[t,t^{-1}]\frac{d}{dt}$
and
$\mathbf{g}(\Gamma)=\left\{g:\Gamma\rightarrow\mathbb{F}[t,t^{-1}]\,|\,(\alpha-\beta)g_{\alpha+\beta}=\alpha g_{\alpha}-\beta g_{\beta},\,\alpha,\beta\in\Gamma\right\},$
where $\mathbb{F}[t,t^{-1}]\frac{d}{dt}$ is a derivation algebra of $\mathbb{F}[t,t^{-1}]$. The following linear maps $D_{\phi},\,D_{g},\,D_{b},\,D^{\rho}$ on $\mathcal{L}(\Gamma)$ are given by, respectively:
$D_{\phi}(L_{\alpha,i})$ $=\phi(\alpha)L_{\alpha,i},$ $D_{\phi}(H_{\alpha,i})$ $=\phi(\alpha)H_{\alpha,i},$
$D_{g}(L_{\alpha,i})$ $=g_{\alpha}H_{\alpha,i},$ $D_{g}(H_{\alpha,i})$ $=0,$
$D_{b}(L_{\alpha,i})$ $=0,$ $D_{b}(H_{\alpha,i})$ $=bH_{\alpha,i},$
$D^{\rho}(L_{\alpha,i})$ $=L_{\alpha}\rho(t^{i}),$ $D^{\rho}(H_{\alpha,i})$ $=H_{\alpha}\rho(t^{i}),\,\mbox{for all}\,\alpha\in\Gamma\,\mbox{and all}\,i\in\mathbb{Z}.$
@@
$\mathcal{D}_{\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}])}\,:=\,\left\{D_{\phi}\,|\,\phi\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}])\right\}$,
@@
$\mathcal{D}_{\mathbb{F}[t,t^{-1}]}\,:=\,\left\{D_{b}\,|\,b\in\mathbb{F}[t,t^{-1}]\right\}$,
$\mathcal{D}_{\mathbb{F}[t,t^{-1}]\frac{d}{dt}}\,:=\,\{D^{\rho}\,|\,\rho\in\mathbb{F}[t,t^{-1}]\frac{d}{dt}\}$
be the corresponding subspaces of $\mathrm{Der}\mathcal{L}(\Gamma)$, respectively. We brief note $\mathbf{g}(\Gamma)$ by $\mathbf{g}$ in the sequel.
@@
Suppose $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$, then there exist $f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}]),\,g\in\mathbf{g}$ such that
$D(L_{\alpha,i})=f_{\alpha}L_{\alpha,i}+g_{\alpha}H_{\alpha,i}$ for all $\alpha\in\Gamma$ and all $i\in\mathbb{Z}.$
@@
Let $D(L_{\alpha,i})=f_{\alpha,i}L_{\alpha,i}+g_{\alpha,i}H_{\alpha,i}$ where $f_{\alpha,i},g_{\alpha,i}\in\mathbb{F}[t,t^{-1}]$. On the one hand, we have
$D([L_{\alpha,i},L_{\beta,j}])=(\alpha-\beta)(f_{\alpha+\beta,i+j}L_{\alpha+\beta,i+j}+g_{\alpha+\beta,i+j}H_{\alpha+\beta,i+j}).$
@@
$D([L_{\alpha,i},L_{\beta,j}])$ == $D(L_{\alpha,i}),L_{\beta,j}]+[L_{\alpha,i},D(L_{\beta,j})]$
== $(\alpha-\beta)f_{\alpha,i}L_{\alpha+\beta,i+j}+\alpha g_{\alpha,i}H_{\alpha+\beta,i+j}+(\alpha-\beta)f_{\beta,j}L_{\alpha+\beta,i+j}$
$\beta g_{\beta,j}H_{\alpha+\beta,i+j}$.
Then we obtain
$f_{\alpha+\beta,i+j}=f_{\alpha,i}+f_{\beta,j},\,\alpha\neq\beta$,
$(\alpha-\beta)g_{\alpha+\beta,i+j}=\alpha g_{\alpha,i}-\beta g_{\beta,j}$ for all $\alpha,\beta\in\Gamma$ and all $i,j\in\mathbb{Z}$.
@@
Thus $f_{0,0}=0$. Let $\beta=-\alpha\neq 0,i=j=0$. Then
$0=f_{0,0}=f_{\alpha,0}+f_{-\alpha,0},\,\alpha\neq 0.$
Let $0\neq\eta\in\Gamma$ satisfying $\eta\neq\alpha,\,-\eta\neq\beta,\alpha+\eta\neq\beta-\alpha$. Then
@@
That is, $f_{\alpha+\beta,i+j}=f_{\alpha,i}+f_{\beta,j}$ for all $\alpha,\beta\in\Gamma$ and all $i,j\in\mathbb{Z}.$ Then $f_{\alpha,i}=f_{\alpha,0}+f_{0,i}$. Let $\rho=tf_{0,1}\frac{d}{dt}$. We have
@@
Replace $D$ by $D-D^{\rho}$. Let $D(L_{0,1})=0$. Then $f_{0,1}=0$ and $f_{0,i}=if_{0,1}=0$, thus $f_{\alpha,i}=f_{\alpha,0}+f_{0,i}=f_{\alpha,0}$ for all $\alpha\in\Gamma$ and all $i\in\mathbb{Z}.$ For convenience, denote $f_{\alpha,i}$ as $f_{\alpha}$. Then
$f_{\alpha+\beta}=f_{\alpha}+f_{\beta},\,f\in\mathrm{Hom}_{\mathbb{Z}}(\Gamma,\mathbb{F}[t,t^{-1}]).$
Let $\alpha=0,j=0$. Then we have
@@
Let $\alpha=-\beta\neq 0$. Then we have
$2\beta g_{0,i+j}$ == $\beta g_{\beta,j}+\beta g_{-\beta,i}$
== $\beta g_{\beta,0}+\beta g_{-\beta,0}$
== $2\beta g_{0,0}$
== 0.
Thus $g_{\beta,i}=g_{\beta,0}-g_{0,i}=g_{\beta,0}$ for all $\beta\in\Gamma$ and all $i\in\mathbb{Z}.$ For convenience, denote $g_{\beta}$ by $g_{\beta,i}$. Then $(\alpha-\beta)g_{\alpha+\beta}=\alpha g_{\alpha}-\beta g_{\beta}$ for all $g\in\mathbf{g}(\Gamma).$
@@
Let $D\in(\mathrm{Der}\mathcal{L}(\Gamma))_{0}$. Then there exists $b\in\mathbb{F}[t,t^{-1}]$ such that $D(H_{\alpha,i})=(f_{\alpha}+b)H_{\alpha,i}$ for all $\alpha\in\Gamma$ and all $i\in\mathbb{Z}.$